what do the equations 3x - 2y = 3 and -3x + 2y = 2 have in common

Answer by stanbon(75887)

what do the equations 3x - 2y = 3 and -3x + 2y = 2 have in common
----
y = (3/2)x-(3/2)
y = (3/2)x + 1
----
The equations have the same slope.
The lines are parallel.
=====================
Cheers,
Stan H.